OpenStudy (anonymous):

Which is the graph of the line with equation: y – 4 = 2(x + 1)

OpenStudy (mathmale):

Why not solve this equation for y first, and then, once you have the slope-intercept form y = mx + b, graph (0,b) and use the slope, m, to construct as many more points on the line as you wish (2 or 3 are enough).

OpenStudy (mathmale):

Another approach: y - 4 = 2(x + 1) could be re-written as y - 4 = 2(x - [-1]). This is the point-slope form of the equation of a straight line. This line goes through (-1,4). Use the slope (2/1) to determine 1 or 2 more points on this line, then draw the line through these points.